Number of postive integral solution of xyz=21600 ?? Please solve this question.....

Grade:12th pass

2 Answers

Arun
6 years ago
we can write 21600 as a product of its prime factors
21600 = 2*2*2*2*2*3*3*3*5*5
21600 = 25 * 33 * 52
hence no. of positive integral solution = (1+2+3+4+5+6)* (1+2+3+4) *(1+2+3)
 = 21 * 10* 6
 = 1260
hope it helps
Joy Gupta
5 years ago
 
 
 
We can use beggers and coins method
21600=6^3*100
21600=2^5*3^3*5^2.         
xyz are beggers so number of solutions=(5+3-1C3-1)(3+3-1C3-1)(2+3-1C3-1)=1260